mr.hart made two investments, one at 4.5% and the other at 3%. the amount invested at 3% exceeded the amount invested at 4.5% by $300. if his total annual income was $84, find the amount invested at each rate.
Answer by nerdybill(7384)

mr.hart made two investments, one at 4.5% and the other at 3%. the amount invested at 3% exceeded the amount invested at 4.5% by $300. if his total annual income was $84, find the amount invested at each rate.
.
Let x = amount ($) invested at 4.5%
then
x+300 = amount ($) invested at 3%
.
.045x + .03(x+300) = 84
.045x + .03x+9 = 84
.075x+9 = 84
.075x = 75
x = $1000 (invested at 4.5%)
.
invested at 3%:
x+300 = 1000+300 = $1300